How many years will it take you to double your money if your rate of return is 7% annually?
Answer #1
FV = PV*(1+r)^n
Substituting values we have,
2 = 1*1.07^n
Taking log of both sides
log2 = n*log1.07
n = log2/log1.07 = 0.30103/0.02938 = 10.25
Answer: Number of years = 10.25
Alternative solution.
Applying rule 72, it will be 72/7 = 10.29
This is a simple approximation.